# Self-organisation - the underlying principle and a general model

### v1

@@ I. Introduction
Most studies of SO have been usually focused on specific contexts and in many case are descriptive. Attempts at finding a general principle underpinning SO are few and they are mainly based on energy considerations. For example, it has been proposed that SO emerges in steady states by selection of microstates that minimise energy-dissipation in driven systems Cyetal21 ; Keetal23 . These models propose that the system random-walks in the multi-dimensional microstates space and those special microstates are selected as the system self-organises. However, this does not explain what is exactly the physical mechanism that makes the random walk park on the microstates of low energy dissipation. Moreover, this idea excludes emergence of SO in out-of-equilibrium a-thermal systems, where energy plays a negligible role. Observations of SO in these systems suggest that a more general principle is in play, which applies to all systems, however driven. Statement of the principle in insufficient - it should also point the way toward a unified modelling approach that leads to prediction of the self-organised state and some relevant properties.
@@ II. SO signatures in 2D granular systems
In steady states of out-of-equilibrium thermal, systems, the invested driving energy is balanced against internal dissipation. The noise in these processes is generated both by the dynamics and, depending on magnitude, also by thermal fluctuations. To illustrate the thesis here, I begin with clarifying the principle in the context of the specific example of the emergence of SO in quasi-static dynamics of driven 2D granular systems. In recent experiments and simulations, the driving methods were simple shear, pure shear, and both isotropic and anisotropic compression. The term quasi-static means that any dynamic process is slower than the response time of stress equilibration, which also means that stopping at any moment, the resulting structural configuration of the assembly of particles is in mechanical equilibrium. It is these configurations that display signatures of SO. Almost all these signatures were seen by studying one structural element - the cells. These are the smallest possible voids in the assembly, which are surrounded by particles in contact. The first signature is that distributions of several structural quantities collapse onto a master forms, once simply scaled by their means. Such a collapse was observed for many (up to 120!) different systems and different packing protocols MaBl14 . Another intriguing signature involves the evolution of cell orders, where an order is defined as the number of particles in contact around a cell. A cell orders change when contacts are made and broken and these processes resemble chemical reactions. It has been found both numerically and experimentally that the dynamics evolve toward steady states that satisfy detailed balance Waetal20 ; Suetal25 . Another observation, providing a clue to the origin of the SO, was of a competition between the stability of cells and their configurational entropy MaBl17 ; Suetal20 . The smoking gun, however, has been observations that, in a number of systems and protocols, there are strong correlations across the systems between the local principal axes of cell shapes and cell stresses MaBl17 ; MaBl21 ; Jietal25 . This strong local alignment demonstrates clearly that the local structure and local stress self-organise cooperatively, a feature that has to be included in any potential model of SO.
@@
These observations offer insight into the principle that underlies the SO in these 2D granular dynamics. The driving and the boundary conditions, generate an internal stress field that fluctuates spatially and temporally. If a cell cannot support the local stress it disintegrates, namely, at least one of the contacts of the particles surrounding it breaks. A system-wide configuration consists of a snapshot of all its cell configurations. The SO phenomenon arises when the occurrence probabilities of a minute subset of cell shapes, orientations, and stresses configurations dominates the observations. The reason for this dominance is that these configurations of cell shapes and stresses are much more stable than the rest and therefore survive much longer. Because the lifetimes of all other configurations are too short they do not feature significantly in a typical state. In the granular systems there is a local coupling between the cell shape and stress, when unstable cells either rearrange or disintegrate, the local stress configuration around them resets. Thus, local stresses persist for longer around more stable cells and stability is correlated with the large principal stress aligning with the cell’s major axis. This interpretation is consistent with all the numerical and experimental observations and, in particular, it is almost surely the reason for the strong system-wide stress-shape orientational correlations.
@@ III. SO statistical mechanics in granular systems
With occurrence probabilities of configurations playing a key role in SO emergence in these dynamics, a natural theoretical model for it is statistical mechanics. To formulate it, several definitions need to be introduced. A cell surrounded by $k$ particles in (force-carrying) contact is a cell of order $k$, or $k$-cell. The cell order distribution across the system is discrete, with $Q_{k}$ ($k=3,4,...,K$) the fractions of $k$-cells out of all $N_{c}$ cells. Cell stability depends on two ’attributes’: its shape and stress. To quantify the former, cells are approximated as ellipses Jietal25 . A cell ellipse has three degrees of freedom (DOFs): its major and minor axes lengths, $a_{c}$ and $b_{c}$, respectively, and the orientation of the major axis relative to a fixed frame of reference, $\theta_{c}$. Cell stresses are defined, using an appropriately weighted sum over the stresses of the particles surrounding it Bletal25 ; SM , as detailed in the supplemental material SM . The cell stress also has three degrees of freedom: the large and small principal stresses, $\sigma_{c1}$ and $\sigma_{c2}$, respectively, and the orientation of the large principal stress relative to the same fixed frame of reference, $\phi_{c}$.
@@
For simplicity, a similar form is assigned to the shape stability,
@@
Next, a survivability function is required, whose value determines the configurations occurrence probabilities. It should include the stability measures, as well as any other constraints on the system. It’s main aim is to quantify the stability, and thence survivability, of system-wide configurations. The observations of function collapses and correlations between stress and shapes further suggest coupling between $\theta_{c}$ and $\phi_{c}$ at the cell level, which the survivability function should include. Additionally, it has been found Jietal25 that the most probable values of the stability measures are finite, which is another piece of information that needs to be accommodated in the survivability function. In principle, it could also include correlations between the DOFs nearest-neighbour cells, although no such correlations have been observed in the experimental and numerical systems Waetal20 ; Suetal25 . The survivability function should decreases as cell stability increases, which would make it analogous to the free energy in thermal systems. Finally, the partition function should also be grand-canonical because the number of cells, $N_{c}$, fluctuates. All these specifications are satisfied by
@@
In this expression, $\mu$, $J_{1}$, and $J_{2}$ are Lagrange multipliers that reflect the effect of the dynamics-generated noise on the number of cells, the cell shape stability, and the cell stress stability, respectively. $f_{0}(k)$ and $h_{0}(k)$ are the most probable values of these measures, found in Jietal25 . The last term represents the cell-cell interaction, with $\psi_{c^{\prime}c}=\left(\theta_{c}-\theta_{c^{\prime}}\right)^{2}$ the difference between the orientations of neighboring cells $c$ and $c^{\prime}$. More interaction terms between other DOFs of neighbouring cells can be added to $F_{c}$, if necessary.
@@
the densities of states provide the occurrence probabilities of the DOFs and/or the cells stability measures. For the 2D granular systems, some of these distributions are known from modelling and observations. Specifically, $\theta_{c}$ is isotropic on the cell level and $\phi_{c}$ is strongly correlated with it Jietal25 . Thus, these two variable can be replaced by one variable, $\Delta\equiv\theta-\phi$. This variable is normally distributed almost identically around zero for all $k$, $P(\Delta)\sim e^{-\Delta^{2}/\left(2\delta^{2}\right)}$. In the absence of other information, the distributions of the stability measures $h$ and $f$ can be obtained from the distributions of the DOFs on which they depend. However, it has been shown that, in these systems, the conditional distributions of $h_{c}$ given $k$, $P_{h}(h|k)$, when scaled by the mean for each cell order, $\bar{h}(k)$, collapse onto an identical Weibull form, $W\left[\hat{h}=h/\bar{h}(k)\right]$, that is independent of $k$ Jietal25 . The observations that no cell-cell spatial correlations exist in two shear experiments Waetal20 ; Suetal25 , simplifies the analysis, $J_{c^{\prime}c}=0$, at least for these systems. This then becomes an ‘independent cells model’, . To illustrate the use of the formalism, assume for simplicity that the conditional densities of states of $f(k)$ also collapse onto a master form under a similar scaling. An explicit calculation of the partition function, detailed in the supplemental material SM , yields
[…]
$-\mu-\frac{1}{2}\ln{\left(1+2\delta^{2}J_{3}\right)}\Bigg{\}}N_{c}\ .$ (5)
This expression can be used to obtain expectation values of quantities, which depend on the DOFs and the stability measures, as functions of the noise parameters and coupling constants in the survivability function,
[…]
@@
The extension of this formalism to three-dimensional (3D) granular SO is straightforward in principle. As in 2D, cells are the smallest polyhedral voids surrounded by particles in contact. They can be approximated as ellipsoids and their stresses defined as weighted sum of the stresses of their surrounding particles. This would give 10 DOFs per cell (three axes and two angles for each of the shape and stress ellipsoids). Similar ellipsoid and stress stability measures can be defined and $F_{c}$ can be a straightforward 3D version of (4). This, in principle, provides the partition function. However, its evaluation may be more difficult than in 3D, especially because the more complicated cells classification into families. In both 2D and 3D, a cell is the smallest void surrounded by particles in contact. This makes for polygons in 2D and polyhedra in 3D. While a polygon can be classified only by its order, $k$, classification of polyhedra requires more parameters: the number of its faces and the order of each polygonal face. Specifically, a family of cell polyhedra is classified by its $n_{p}$ vertices (contacts), $k_{3}$ triangular faces, $k_{4}$ rectangular faces, and so on up to the highest-order faces, $k_{t}$. Members of the same family have an identical series of such numbers. However, this is a quantitative, rather than qualitative, complication. It just means that the product over $k$ in (3) needs to be replaced by several products, while the integrand remains the same. This extension is under investigation currently, but it is downstream from the aim here and will be presented elsewhere.
@@ IV. The principle and its generalisation
This principle can be extended beyond granular systems. Driving a disordered system generates noise even if thermal fluctuations are negligible. Such systems are also constrained by the environment through boundary conditions and interactions with other dynamic systems, as well as internally by the nature of interactions between their individual components. SO emerges when only a minute subset of system configurations, its microstates, survive the noise sufficiently long to dominate experimental observations. In contrast to these, the bulk of system-wide configurations disintegrate faster than observational timescales and therefore are not typical. Different driving methods may give rise to SOs of different characteristics because it is the noise that determines which configurations survive and the noise is generated by the driving.
@@ V. Extension of the modelling framework
The extension to more general modelling of SO is as follows. (1) define quasi-particles (the cells); (2) identify the quasi-particles’ DOFs (the cell ellipse axes lengths and orientations, principal stresses and orientation); (3) identify quasi-particle stability measures that depend on the DOFs and whose values decrease with increasing quasi-particle stability (the cell stresses and shapes); (4) construct a survivability function that depends on all the quasi-particles’ stability measures and decreases monotonically as their stability increases ($F_{c}$ in eq. (4)). The value of the survivability function at each system-wide configuration is determined by the configuration’s stability, which, in turn, determines the configuration’s occurrence probability and therefore its lifetime.
Consider then a system of $N$ elements, structural or otherwise, which show signatures of SO, and define these as quasi-particles, $q=1,2,...,N$. Associated with each, are $M$ properties, indexed $m=1,2,...,M$. These properties self-organise locally for every quasi-particle across the system. Property $m$ of quasi-particle $q$ depends on $A_{m}$ of its DOFs, indexed $a_{q1}^{m},a_{q2}^{m},...,a_{qA_{m}}^{m}$. The modelling of this system requires identification of $f^{m}_{q}$ stability measure of each property $m$ of quasi-particle $q$. The quasi-particle survivability function, $F_{q}$, depends on these measure, as well as on other DOFs. It can also involve interactions between DOFs of the same quasi-particles, as well as different ones. For example, a generalisation of (4) would be
@@ VI. Example application of the formalism
To illustrate the application of this formalism, consider the phenomenon of self-organised laning He71 ; HeMo95 ; Chetal00 . Laning refers to the spontaneous formation of lanes as agents (pedestrians, cars, active agents, etc.) move in a relatively crowded spaces. It is the result of agents aiming to move in a particular direction or reach a particular destination without bumping too often and too violently into others. While collisions, and their avoidance, are local, ordered system-wide laning often emerges. This SO makes the flow of agents more efficient, e.g., by minimising the time spent in the crowd before reaching the destination. In a simple version of this problem, $N$ agents move in one direction along a long strip of width that allows only few agents to move side by side. They can move at one of two speeds $v_{1}$ and $v_{2}>v_{1}$ and at any angle relative to the forward direction, $-\pi/2<\theta<\pi/2$. Moving at $\theta=0$ is fastest, but can be hampered by collisions with more slowly-moving agents. Attempting to overtake those by moving sideways may incur collisions with other agents. Here, the agents are the quasi-particles and $q=1,2,...,N$ and each agent has two DOFs,the speed and direction of movement. For each agent, the most stable mode of movement is by avoiding collisions, achievable with the stability measure,
@@
where $q^{\prime}\Rightarrow q$ denotes the indices of the agents nearest to $q$ (identifiable, e.g., using Voronoi tessellation). The objective of reaching the destination efficiently is a constraint that also needs to be quantified in the survivability function and this is achievable by encouraging movement in the forward direction. Thus, a natural survivability function is
@@
×$\times\prod\limits_{q=1}^{N}\Big{\{}P\left(v_{q},\theta_{q}\right)dv_{q}d\theta_{q}\Big{\}}\ ,$ (13)
with, say, $P\left(v_{q},\theta_{q}\right)=\left[p\delta(v-v_{1})+(1-p)\delta(v-v_{2})\right]/\pi$ the density of states. It is clear that, by increasing $J$ and $\lambda$, the stable system-wide configurations are in a small subset that approximate two lanes, one moving at $v_{1}$ and the other at $v_{2}$. This partition function can be readily evaluated, but this is downstream from the aim of this example, which is only to illustrate the application of the general formalism.

### v2

@@ I. Introduction
Most studies of SO have been usually focused on specific contexts and in many case are descriptive. Attempts at finding a general principle underpinning SO are few and they are mainly based on energy considerations. For example, it has been proposed that SO emerges in steady states by selection of microstates that minimise energy-dissipation in driven systems Cyetal21 ; Keetal23 . These models propose that the system random-walks in the multi-dimensional microstates space and those special microstates are selected as the system self-organises. However, this does not explain what is exactly the physical mechanism that makes the random walk park on the microstates of low energy dissipation. Moreover, this idea excludes emergence of SO in out-of-equilibrium a-thermal systems, where energy plays a negligible role. Observations of SO in these systems suggest that a more general principle is in play, which applies to all systems, however driven. Statement of the principle is insufficient - it should also point the way toward a unified modelling approach that leads to prediction of the self-organised state and some relevant properties.
@@ II. SO signatures in 2D granular systems
In steady states of out-of-equilibrium thermal systems, the invested driving energy is balanced against internal dissipation. The noise in these processes is generated both by the dynamics and, depending on magnitude, also by thermal fluctuations. To illustrate the thesis here, I begin with clarifying the principle in the context of the specific example - the emergence of SO in quasi-static dynamics of driven 2D granular systems. In recent experiments and simulations, the driving methods were simple shear, pure shear, and both isotropic and anisotropic compression. The term quasi-static means that any dynamic process is slower than the response time of stress equilibration, which also means that stopping at any moment, the resulting structural configuration of the assembly of particles is in mechanical equilibrium. It is these configurations that display signatures of SO. Almost all these signatures were seen by studying one structural element - the cells. These are the smallest possible voids in the assembly, which are surrounded by particles in contact. The first signature is that distributions of several structural quantities collapse onto a master forms, once simply scaled by their means. Such a collapse was observed for many (up to 120!) different systems and different packing protocols MaBl14 . Another intriguing signature involves the evolution of cell orders, where an order is defined as the number of particles in contact around a cell. Cell orders change when contacts are made and broken and these processes resemble chemical reactions. It has been found both numerically and experimentally that the dynamics evolve toward steady states that satisfy detailed balance Waetal20 ; Suetal25 . Another observation, providing a clue to the origin of the SO, was of a competition between the stability of cells and their configurational entropy MaBl17 ; Suetal20 . The smoking gun, however, has been observations that, in a number of systems and protocols, there are strong correlations across the systems between the local principal axes of cell shapes and cell stresses MaBl17 ; MaBl21 ; Jietal25 . This strong local alignment demonstrates clearly that the local structure and local stress self-organise cooperatively, a feature that has to be included in any potential model of SO.
@@
These observations offer insight into the principle that underlies the SO in these 2D granular dynamics. The driving and the boundary conditions generate an internal stress field that fluctuates spatially and temporally. If a cell cannot support the local stress it disintegrates, namely, at least one of the contacts of the particles surrounding it breaks. A system-wide configuration consists of a snapshot of all its cell configurations. The SO phenomenon arises when the occurrence probabilities of a minute subset of cell shapes, orientations, and stresses configurations dominates the observations. The reason for this dominance is that these configurations of cell shapes and stresses are much more stable than the rest and therefore survive much longer. All other configurations are not observed in simulations and experiments because their lifetimes are too short. In the granular systems, the cell shape and cell stress are locally coupled. When unstable cells either rearrange or disintegrate, the stress around them also resets. Thus, local stresses persist for longer around more stable cells and stability is correlated with the large principal stress aligning with the cell’s major axis. This interpretation is consistent with all the numerical and experimental observations and, in particular, it is almost surely the reason for the strong system-wide stress-shape orientational correlations.
@@ III. SO statistical mechanics in granular systems
With occurrence probabilities of configurations playing a key role in SO emergence in these dynamics, a natural theoretical model for it is statistical mechanics. To formulate it, several definitions need to be introduced. A cell surrounded by $k$ particles in (force-carrying) contact is a cell of order $k$, or $k$-cell. The cell order distribution across the system is discrete, with $Q_{k}$ (k=3,4,,Kk=3,4,...,K) the fractions of $k$-cells out of all $N_{c}$ cells. Cell stability depends on two ’attributes’: its shape and its stress. To quantify the former, cells are approximated as ellipses Jietal25 . A cell ellipse has three degrees of freedom (DOFs): its major and minor axes lengths, $a_{c}$ and $b_{c}$, respectively, and the orientation of the major axis relative to a fixed frame of reference, $\theta_{c}$. Cell stresses are defined, using an appropriately weighted sum over the stresses of the particles surrounding it Bletal25 , as detailed in the supplemental material SM . The cell stress also has three degrees of freedom: the large and small principal stresses, $\sigma_{c1}$ and $\sigma_{c2}$, respectively, and the orientation of the large principal stress relative to the same fixed frame of reference, $\phi_{c}$.
@@
Any model for the shape stability can be used and, for illustration purposes, a similar form is assigned here,
@@
Next, a survivability function is required, whose value determines the configurations occurrence probabilities. It should include the stability measures, as well as any other constraints on the system. It’s main aim is to quantify the stability, and thence survivability, of system-wide configurations. The observations of function collapses and correlations between stress and shapes further suggest coupling between $\theta_{c}$ and $\phi_{c}$ at the cell level, which the survivability function should include. Additionally, it has been found Jietal25 that the most probable values of the stability measures are finite, which is another piece of information that needs to be accommodated in the survivability function. In principle, it could also include correlations between the DOFs of nearest-neighbour cells, although no such correlations have been observed in the experimental and numerical systems Waetal20 ; Suetal25 . The survivability function should decreases as cell stability increases, which would make it analogous to the free energy in thermal systems. Finally, the partition function should also be grand-canonical because the number of cells, $N_{c}$, fluctuates. All these specifications are satisfied by
@@
In this expression, $\mu$, $J_{1}$, and $J_{2}$ are Lagrange multipliers that reflect the effect of the dynamics-generated noise on the number of cells, the cell shape stability, and the cell stress stability, respectively. The last term represents the cell-cell interaction, with $\psi_{c^{\prime}c}=\left(\theta_{c}-\theta_{c^{\prime}}\right)^{2}$ the difference between the orientations of neighboring cells $c$ and $c^{\prime}$. More interaction terms between other DOFs of neighbouring cells can be added to $F_{c}$, if necessary.
@@
the densities of states provide the occurrence probabilities of the DOFs and/or the cells stability measures. For the 2D granular systems, some of these distributions are known from modelling and observations. Specifically, $\theta_{c}$ is isotropic on the cell level and $\phi_{c}$ is strongly correlated with it Jietal25 . Thus, these two variable can be replaced by one variable, $\Delta\equiv\theta-\phi$. This variable is normally distributed almost identically around zero for all $k$, $P(\Delta)\sim e^{-\Delta^{2}/\left(2\delta^{2}\right)}$. In the absence of other information, the distributions of the stability measures $h$ and $f$ can be obtained from the distributions of the DOFs on which they depend. However, it has been shown that, in these systems, the conditional distributions of $h_{c}$ given $k$, $P_{h}(h|k)$, when scaled by the mean for each cell order, $\bar{h}(k)$, collapse onto an identical Weibull form, $W\left[\hat{h}=h/\bar{h}(k)\right]$, that is independent of $k$ Jietal25 . Observations suggest that, at least in two shear experiments, there are no cell-cell spatial correlations Waetal20 ; Suetal25 . Namely, $J_{c^{\prime}c}=0$ and this becomes an ‘independent cells model’. To demonstrate the use of the formalism, assume for simplicity that the conditional densities of states of $f(k)$ also collapse onto a master form under a similar scaling. An explicit calculation of the partition function, detailed in the supplemental material SM , yields
$\ln{Z}=\Bigg{\{}$ $\sum\limits_{k=1}^{K}Q_{k}\ln\left[\langle e^{-J_{1}\bar{f}(k)}\rangle\langle e^{-J_{2}\bar{h}(k)}\rangle\right]$
$-\mu-\frac{1}{2}\ln{\left(1+2\delta^{2}J_{3}\right)}\Bigg{\}}N_{c}\ ,$ (5)
with the angular brackets denoting averages over all the DOFs. This expression can be used to obtain expectation values of quantities, which depend on the DOFs and the stability measures, as functions of the noise parameters and coupling constants in the survivability function,
$\langle G\rangle=\frac{1}{Z}\Bigg{\{}$ $\prod\limits_{k=1}^{K}e^{-\mu}\int G\left(\{X\}\right)e^{-J_{1}f-J_{2}h-J_{3}(\theta-\phi)^{2}}$
@@
The extension of this formalism to three-dimensional (3D) granular SO is straightforward in principle. As in 2D, cells are the smallest polyhedral voids surrounded by particles in contact. They can be approximated as ellipsoids and their stresses defined as weighted sum of the stresses of their surrounding particles. This would give 10 DOFs per cell (three axes and two angles for each of the shape and stress ellipsoids). Similar ellipsoid and stress stability measures can be defined and $F_{c}$ can be a straightforward 3D version of (4). This, in principle, provides the partition function. However, its evaluation may be more difficult than in 3D, especially because of the more complicated cells classification into families. Cells are polygons in 2D and polyhedra in 3D and, while a polygon can be classified only by its order, classification of polyhedra requires more parameters: the number of its faces and the order of each polygonal face. Specifically, a family of cell polyhedra is classified by its $n_{p}$ vertices (contacts), $k_{3}$ triangular faces, $k_{4}$ rectangular faces, and so on up to the highest-order faces, $k_{t}$. Members of the same family have an identical series of such numbers. Nevertheless, this complication is a quantitative, rather than qualitative. It just means that the product over $k$ in (3) needs to be replaced by products over all the cell parameters, but the integrand remains the same. This extension is under investigation currently, but it is downstream from the aim here and will be presented elsewhere.
@@ IV. The principle and its generalisation
This principle can be extended beyond granular systems. Driving a disordered system generates noise even if thermal fluctuations are negligible. Such systems are also constrained by the environment through boundary conditions and interactions with other dynamic systems, as well as internally by the nature of interactions between their individual components. SO emerges when only a minute subset of system configurations, its microstates, survive the noise sufficiently long to dominate experimental observations. In contrast to these, the bulk of system-wide configurations disintegrate faster than observational timescales and therefore are not typical. As it is the driving that determines the noise and it is the noise that determines which configurations survive, different driving methods may give rise to SOs of different characteristics.
@@ V. Extension of the modelling framework
The extension to more general modelling of SO is as follows. (1) define quasi-particles (e.g., the cells); (2) identify the quasi-particles’ DOFs (the cell ellipse axes lengths and orientations, principal stresses and orientation); (3) identify quasi-particle stability measures that depend on the DOFs and whose values decrease with increasing quasi-particle stability (the cell stresses and shapes); (4) construct a survivability function that depends on all the quasi-particles’ stability measures and decreases monotonically as their stability increases ($F_{c}$ in eq. (4)). The value of the survivability function at each system-wide configuration is determined by the configuration’s stability, which, in turn, determines the configuration’s occurrence probability and therefore its lifetime.
Consider then a system of $N$ elements, structural or otherwise, which show signatures of SO, and define these as quasi-particles, q=1,2,,Nq=1,2,...,N. Associated with each, are $M$ properties, indexed m=1,2,,Mm=1,2,...,M. These properties self-organise locally for every quasi-particle across the system. Property $m$ of quasi-particle $q$ depends on $A_{m}$ of its DOFs, indexed $a_{q1}^{m},a_{q2}^{m},...,a_{qA_{m}}^{m}$. The modelling of this system requires identification of $f^{m}_{q}$ stability measure of each property $m$ of quasi-particle $q$. The quasi-particle survivability function, $F_{q}$, depends on these measure, as well as on other DOFs. It can also involve interactions between DOFs of the same quasi-particles, as well as different ones. For example, a generalisation of (4) would be
@@ VI. Example application of the formalism
To illustrate the application of this formalism, consider the phenomenon of self-organised laning He71 ; HeMo95 ; Chetal00 . Laning is a spontaneous formation of lanes as agents (pedestrians, cars, etc.) move in crowded spaces. It is the result of agents aiming to move in a particular direction or reach a particular destination without bumping too often and too violently into others. While collisions, and their avoidance, are local, ordered system-wide laning often emerges. This SO makes the flow of agents more efficient, e.g., by minimising the time spent in the crowd before reaching the destination. The dynamic lanes pattern is but a minute fraction of all the possible ways that the agents can move. In a simple version of this problem, $N$ agents move in one direction along a long strip of width that allows only few agents to move side by side. They can move at one of two speeds $v_{1}$ and $v_{2}>v_{1}$ and at any angle relative to the forward direction, $\pi/2<\theta<\pi/2$. Moving at $\theta=0$ is fastest, but can be hampered by collisions with more slowly-moving agents. Attempting to overtake those by moving sideways may incur collisions with other agents. Here, the agents are the quasi-particles and q=1,2,,Nq=1,2,...,N and each agent has two DOFs,the speed and direction of movement. For each agent, the most stable mode of movement is by avoiding collisions, achievable with the stability measure,
@@
where $q^{\prime}\Rightarrow q$ denotes the indices of the agents nearest to $q$ (identifiable, e.g., using Voronoi tessellation at each moment). The objective of reaching the destination efficiently is a constraint that also needs to be quantified in the survivability function and this is achievable by encouraging movement in the forward direction. Thus, a natural survivability function is
@@
$\times\prod\limits_{q=1}^{N}\Big{\{}P\left(v_{q},\theta_{q}\right)dv_{q}d\theta_{q}\Big{\}}\ ,$ (13)
with, say, $P\left(v_{q},\theta_{q}\right)=\left[p\delta(v-v_{1})+(1-p)\delta(v-v_{2})\right]/\pi$ the density of states. It is clear that, by increasing $J$ and $\lambda$, the stable system-wide configurations are in a small subset that approximate two lanes, one moving at $v_{1}$ and the other at $v_{2}$. This partition function can be readily evaluated, but, as the aim here is only to illustrate the application of the general formalism, this is downstream from the current work.